OpenStudy (anonymous):

What is the standard deviation of the following data set rounded to the nearest tenth? 7.7, 8.4, 9, 8, 6.9 a.0.4 b.0.7 c.0.5 d.0.8

OpenStudy (goldphenoix):

It takes a while to do the standard deviation. To find the standard deviation, you want to first find the mean of the data set. So what is the mean?

OpenStudy (anonymous):

8.2

OpenStudy (goldphenoix):

Are you sure you did it right? I got 8.

OpenStudy (anonymous):

nevermind i got 8 now

OpenStudy (goldphenoix):

Alright. Next, you want to subtract the mean from each data. So: 7.7 - 8 = ? 8.4 - 8 = ? 9 - 8 = ? 8 -8 = ? 6.9 - 8 =? Fill in the question mark please.

OpenStudy (anonymous):

-.3 .4 1 0 -1.1

OpenStudy (goldphenoix):

Alright. Next, you want to square them. -0.3^2 = .4^2 = 1^2 = 0^2 = -1.1^2 =

OpenStudy (anonymous):

.09 .16 1 0 1.21

OpenStudy (goldphenoix):

Next, you want to find the mean of it. What is the mean of it?

OpenStudy (anonymous):

.492 or .5

OpenStudy (goldphenoix):

It's .492. Now, .492 is the variance. To find the standard deviation, you want to square root it. So what's the square root of .492?

OpenStudy (anonymous):

.7

OpenStudy (goldphenoix):

Yes. So the standard deviation of that data set it 0.7. I hope this help! Also, http://www.mathsisfun.com/data/standard-deviation.html will help you solve these questions.

OpenStudy (anonymous):

awesome